OpenStudy (anonymous):

A roller coaster starts at rest from a height of 110 m, and accelerates down the track to a height of 10.0 m. Find the speed it can reach, assuming no friction.

OpenStudy (anonymous):

110-10=.5*9.8*t^2 solve for t then v=t*a I think

OpenStudy (ujjwal):

\[v ^{2}=2gh\] since initial velocity is zero here h=110-10 find v

OpenStudy (anonymous):

here only conservative forces act and thus mechanical energy is conserved. \[\Delta PE = KE \] \[mg \Delta h = 1/2 mv^2\] \[v=\sqrt{2gh}\]